BONUS The time it takes a worker on an assembly line to complete a task is exponentially distributed with a mean of 7.8 minutes. What is the probability that it will take a worker less than 5 minutes to complete the task? What is the probability that it will take a worker between 4 and 11 minutes to complete the task? What is the probability that it will take a worker greater than 7.5 minutes to complete the task? a. b. c. 3,4, 3 m
Answer #1

here this is exponential distriution with mean \beta =7.8

a)

P(it takes less than 5 minutes)=P(X<5)=1-e-x/\beta =1-e-5/7.8 =0.4732

b)

P(4<X<11)=P(X<11)-P(X<4)=(1-e-11/7.8 )-(1-e-4/7.8)=0.3547

c)

P(X>7.5)=1-P(X<7.5)=1-(1-e-7.5/7.8)=0.3823